An outside loudspeaker (considered a small source) emits sound waves with a power output of 111 w (a) Find the intensity 10.7 m from the source W/m2 (b) Find the intensity level in decibels at that distance. dB (c) At what distance would you experience the sound at the threshold of pain, 120 dB? Need Help? Read 11. -/2 points SerCP10 14 P023 My No A commuter train passes a passenger platform at a constant speed of 42.1 m/s. The train horn is sounded at its characteristic frequency of 330 Hz. (a) What overall change in frequency is detected by a person on the platform as the train moves from approaching to receding? (Indicate whether increases of decreases with the sign of your answer.) Hz (b) What wavelength is detected by a person on the platform as the train approaches?
Answer #1

10) intensity = P/A = 111/4pi*10.7^2

= 0.76 W/m^2

b) dB = 10 log ( I /Io)

dB = 10 log ( 0.76 / 10^-12)

dB = 118.8 dB

c) dB = 10 log(I/Io)

120 = 10 log(I/10^-12)

I = 1 W/m^2

Area = P/I

4pi*r^2 = 111/1

4pi*r^2 = 111

r = 2.97 m